OpenStudy (anonymous):

During the flu season, 1/5 of the total population is infected. A big school decides to examine its students for flu infection. What is the probability that when the students are examined one at a time, the first student to have the flu is the fourth student examined? Hint: Think of geometric random variables.

OpenStudy (dumbcow):

http://en.wikipedia.org/wiki/Geometric_distribution p = 1/5 \[P(k=4) = (\frac{4}{5})^{3} \frac{1}{5}\]

OpenStudy (anonymous):

Can you explain step by step how you did this?

OpenStudy (kropot72):

The explanation at the link given by @dumbcow looks very good. Please look at it carefully @anneflys.